Fermi National Laboratory
 
Operation Summary for January 4, 2012 - January 6, 2012
 

Wednesday January 4

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI) , and with the NuMI and MiniBooNE experiments taking beam .

During the day shift, a Muon Rings person completed a study at 8 AM. At 8:33 AM, three Booster RF stations (BRF4, BRF10, & BRF14) tripped off multiple times over the next seven hours; they all reset. Operators dumped the Recycler beam and injected new protons at 11:46 AM. A Muon Rings person began a high intensity study at 2:53 PM. Operations received permission to send beam to a new Meson FTBF experiment T-958 at 3:34 PM.

During the evening shift, Operations established beam to T-958 at 4:48 PM. Operators tuned the 750 KeV Line at 6 PM. Operations disabled T-958 beam at 10:05 PM. At 11:07 PM, the MiniBooNE beam permit tripped off due to a Switch magnet’s (MBEX) power supply drifting low; Operators power cycled the supply.

Thursday January 5

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, at 4:33 AM, the I- Source sparked and its intensity dropped; Operators began its recovery. Operations turned off MI and the Booster at 6:30 AM in preparation for accesses. A Preaccelerator person took down the I- Source for repairs at 6:59 AM. Operators reported that by the end of shift, the following areas were in access: Booster, MI, F-Sector, Transfer Hall, the Muon Rings Pre-Vault and Pre-Target, the MI-60 South power supply kicker room, and the MI-dump enclosure.

During the day shift, Operators prepared NuMI for access at 8:56 AM. At 9:27 AM, personnel working in the MI tunnel noticed damaged vacuum pump power supply cable; since this cable must be repaired the OpsCo extended the shutdown. Operators reported at 10:33 AM that the Transfer Hall, F-Sector, and MI-65 accesses had ended. The Booster and Muon Rings were out of access by 11:21 AM. Operators reported at 11:28 AM that Power Supply personnel had completed their work on MBEX. Operations established beam to MiniBooNE at 11:33 AM. Operations established beam to NuMI at 11:42 AM. Operations established beam to T-958 at 1:28 PM. At 1:35 PM, Linac personnel reported that during their access they had switched over to the H- Source and had been tuning since then.

During the evening shift, Operations experienced some Booster RF trips, all trips were reset. Operations also experienced a few NuMI permit trips, all were due to an apparent drop in current on LAM60. Operations disabled T-958 beam at 10:09 PM.

Friday January 6

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE .

During the midnight shift, at 12:04 AM, Operations experienced lots of Booster losses due to missing 15HZ resets, Operations took the Master switch (disabling all beam) while they investigate. Beam returned at 12:49 AM when Operations found out the McClure box (TTL fanout) had put itself into MDAT mode, they changed it back to TCLK mode and beam returned. Operations turned on Mubar rings, transport, pre-target and pre-vault at 6:22 AM in preparation for high intensity studies.

Operations Coordinator (OpsCo) Plans for Today and the Weekend

The OpsCo approved of the plans by Booster personnel to conduct studies . We should be ready for a smooth weekend of Operation.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/06/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory